notspam84:

The polynomial x2 + 3x - 40 represents the box’s area. Find the length and the width of the box written as two binomials. Please help and explain

jhonyy9:

so you need factorizing completely this quadratic try get two numbers with sum equal +3 and product equal -40

notspam84:

ive figured out the answer which is (x-5)(x+8), but i dont know how to explain it can you help with that?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so you need factorizing completely this quadratic try get two numbers with sum equal +3 and product equal -40 \(\color{#0cbb34}{\text{End of Quote}}\) look what i ve said above you get (x-5)(x+8) 8+(-5) = +3 8*(-5) = -40 right ?

notspam84:

yes

jhonyy9:

@supie opinion pls ? ty

jhonyy9:

so what result from this ? what are the length and width of the box ?

notspam84:

idk it only gave the area

jhonyy9:

yes this is the area x^2 +3x -40 = (x-5)(x+8) yes ?

notspam84:

yes, but i have to explain how i did that but i dont know the algebra words to do it

notspam84:

like do we use the greatest common factor to find this answer?

notspam84:

or not

notspam84:

please?

jhonyy9:

no you just need rewrite this quadratic binomial

jhonyy9:

@Laylalyssa can you explain it more easy understanding ? ty

notspam84:

Can you explain how you found the answer

jhonyy9:

x^2 +3x -40 = x*x +8x -5x +8*(-5) = x(x+8) -5(x+8) = (x-5)(x+8)